OpenStudy (anonymous):

in 1980, median family income was about $17,000 and in 2000 it was about $42,000. A) find the slope of the line passing through the points (1980,17000) and (2000,42000). B) interpret the slope as a rate of change. C) if this trend continues, estimate the median family income in 2004.

OpenStudy (jazzyfa30):

@Jhannybean help please

jimthompson5910 (jim_thompson5910):

for part a), use the slope formula m = (y2-y1)/(x2-x1)

jimthompson5910 (jim_thompson5910):

In this case, (x1, y1) = (1980, 17000), which means x1=1980 and y1=17000 and (x2, y2) = (2000, 42000), telling us that x2=2000 and y2=42000

OpenStudy (jhannybean):

Let's see. slope is defined by \[m=\] ....i think jim's got this xD

OpenStudy (anonymous):

so it would be m= 42000 - 17000 2000 - 1980

jimthompson5910 (jim_thompson5910):

keep going

OpenStudy (anonymous):

which becomes 25000 20

jimthompson5910 (jim_thompson5910):

good, whats next

OpenStudy (anonymous):

dont know now this is where I kept getting stuck at

jimthompson5910 (jim_thompson5910):

well you can cancel off a 0 from each to get 2500/2

jimthompson5910 (jim_thompson5910):

now either use a calculator or long division to get 2500/2 = ???

OpenStudy (anonymous):

which will give me 1250

jimthompson5910 (jim_thompson5910):

so that's your slope

jimthompson5910 (jim_thompson5910):

what does it mean in this context here?

OpenStudy (anonymous):

the first part does ask what m equals but now it asks if the income increases or decreases and if so if it does by year or between the years of 1980 and 2000

jimthompson5910 (jim_thompson5910):

well it's $17,000 in the year 1980 and it turns into $42,000 in the year 2000

jimthompson5910 (jim_thompson5910):

so you tell me

OpenStudy (anonymous):

I know that it is increasing but I am not sure of yearly or just between them years

jimthompson5910 (jim_thompson5910):

do you think the increase jumps from 17,000 to 42,000 in one big jump?

OpenStudy (anonymous):

yearly

jimthompson5910 (jim_thompson5910):

that's more realistic

jimthompson5910 (jim_thompson5910):

and that's exactly what the slope is: the yearly change in median family income

jimthompson5910 (jim_thompson5910):

m = 1250 means the median family income increases $1250 each year (roughly)

OpenStudy (anonymous):

ok so now to figure out the income for 2004 I would just keep adding 1250 to the 42000 correct

jimthompson5910 (jim_thompson5910):

you could

jimthompson5910 (jim_thompson5910):

that's the long way to do it

OpenStudy (anonymous):

1250 · 4

jimthompson5910 (jim_thompson5910):

and it's kinda error prone if you don't keep track of everything

jimthompson5910 (jim_thompson5910):

or you could find the equation of the line that goes through the two points first

jimthompson5910 (jim_thompson5910):

then plug in x = 2004 to find y

jimthompson5910 (jim_thompson5910):

oh i see what you mean, yeah that works

OpenStudy (anonymous):

in 2004 the income will be 47000 am I correct?

jimthompson5910 (jim_thompson5910):

4*1250 = 5000 so adding 5000 to the income in 2000 will give you the income in 2004

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

ok thank you so much

jimthompson5910 (jim_thompson5910):

yw